OpenStudy (anonymous):

Part A: Write the expression x^2-18x+45 as a product of two linear expressions. Show your work and justify each step. Part B: Rewrite x^2 6x+9 as a square of a linear expression. Part C: Do the expression in parts A and B have a common factor? Justify your answer.

OpenStudy (anonymous):

Use the AC factoring method: x^2 - 18x + 45 = (x - 3)(x - 5) x^2 - 6x + 9 = (x -3)^2 They have a common factor (x -3)